### v1

@@ Abstract
Understanding the dynamics of magnons in quantum magnetism is a captivating area of condensed matter physics. Magnons exhibit unique collective behaviours, including the formation of multimagnon bound states (MBSs). Predicted by Bethe over 90 years ago, MBSs in one-dimensional (1D) quantum magnets remain a significant subject of study due to their fundamental and technological implications. While recent advancements in quantum simulation with ultracold atoms have enabled the observation of MBSs in optical lattices, observing these states in naturally occurring materials presents an even more exciting prospect.
This study proposes a novel mechanism for the stabilisation of MBSs in quasi-1D (Q1D) edge-shared cuprates, materials known for their complex magnetic interactions. Contrary to conventional approaches that require high external magnetic fields, our mechanism leverages small antiferromagnetic (AFM) interchain couplings, which act analogously to an internal magnetic field. These couplings induce collinear AFM order within individual chains, facilitating the formation of stable MBSs. This self-organised magnon condensation driven by interchain interactions is experimentally more accessible and opens new avenues for observations.
Our findings offer a fresh perspective on magnetic ordered states driven by interchain-coupling-induced internal magnetic fields. We provide a comprehensive theoretical framework, present numerical results compared with experimental data from specific Q1D cuprates like Li2CuO2, Ca2Y2Cu5O10, LiCuSbO4, and PbCuSO4(OH)2, and discuss the broader implications of our discoveries. This work paves the way for future research exploring the potential of interchain coupling-driven magnon condensation, significantly advancing the fields of quantum computing, spintronics, and magnonics.
@@
magnon bound state, internal field, quasi-one-dimensional cuprates, density-matrix renormalization group
@@ 1 Introduction
Understanding the dynamics of magnons in quantum magnetism is a captivating area of condensed matter physics. Magnons, the quasiparticles associated with spin waves in magnetically ordered systems, exhibit unique collective behaviours, including the formation of multimagnon bound states (MBSs). The existence of MBSs in one-dimensional (1D) quantum magnets was predicted over 90 years ago by Bethe [1]: Due to ferromagnetic (FM) interactions, two spin excitations can remain bound together, forming what is known as a two-magnon bound state [1, 2, 3]. Investigating the formation and behaviour of MBSs addresses fundamental questions in physics and holds significant potential for advancing quantum computing [4], spintronics [5], and other next-generation technologies like magnonics [6].
@@
Recent advancements in quantum simulation with ultracold atoms have enabled the observation of magnon excitations and their two-magnon bound states in quantum Heisenberg chains comprising bosonic atoms in optical lattice experiments [7, 8]. Nonetheless, observing MBSs in naturally occurring materials would be even more exciting. Pursuing this possibility, theoretical studies have extensively investigated the emergence of MBSs in frustrated ferromagnetic-antiferromagnetic (FM-AFM) $J_{1}$-$J_{2}$ chain systems, which are typical magnetic models for quasi-1D (Q1D) edge-shared cuprates [9]. These studies have demonstrated that MBSs can be induced by a magnetic field, and interestingly, near saturation, the types of MBSs vary with the ratio of $J_{1}$ to $J_{2}$, transitioning between two-magnon (nematic), three-magnon (triatic), four-magnon (quartic), and so on [10, 11, 12, 13, 14, 15, 16, 17]. Subsequent experimental efforts have aimed to identify MBSs, focusing on highly 1D materials such as LiVCuO4, where signatures of nematic states have been observed [18, 19, 20, 21, 22, 23]. Nevertheless, the stabilisation of MBSs generally requires high external magnetic fields, making experimental observation challenging due to their fragility against thermal fluctuations and external perturbations [24, 25].
In this study, we propose a novel mechanism for the stabilisation of MBSs in Q1D edge-shared cuprates, a class of materials known for their complex magnetic interactions. Unlike conventional approaches, our mechanism does not rely on external magnetic fields for the emergence of MBSs. Instead, small AFM interchain couplings act analogously to an internal magnetic field. These interchain couplings induce a collinear AFM order (CAFO) within individual chains, leading to the formation of stable MBSs. This self-organised condensation of magnons driven by interchain interactions is much more accessible for experimental verification, opening in this way new avenues for observations. Although this perspective has been theoretically suggested, it has been somewhat overlooked when considering real materials. Thus, our findings provide a fresh perspective on revisiting various magnetically ordered states as driven by interchain-coupling-induced effective magnetic fields – we call it internal fields.
This paper is structured as follows: we first outline the theoretical framework for MBS and the role of AFM interchain couplings. Next, we present our numerical results and their comparison with experimental data from specific Q1D cuprates. Finally, we discuss the broader implications of our findings and propose future research directions to further explore the potential of interchain coupling-driven magnon condensation.
@@ 2 Magnon bound states by external field
Figure 1: (a) Schematic illustration of the MBS driven by interchain coupling, corresponding to a triatic state with three magnons bound together. The yellow lines represent transverse spin fluctuations. Crystal structures of (b) Li2CuO2, showing two CuO2 chains per unit cell along the b axis, and (c) Ca2Y2Cu5555O10, showing two CuO2 chains per unit cell along the a axis. Red and blue arrows in (b) and (c) indicate an example of symmetry-broken magnetization directions in the CAFO state. Structure of interchain couplings for compounds such as (d) Li2CuO2 and Ca2Y2Cu5555O10, and (e) LiCuSbO4 and PbCuSO4(OH)2. (f) Schematic of the chain positions used in the numerical 2-chain, 4-chain, and 8-chain calculations, where the dashed lines denote the interchain coupling (also, see Ref. [26]).
Prior to discussing our novel emergence mechanism of MBSs, it is necessary to outline the concept of MBSs and to explain how to theoretically calculate the magnon binding energy, which is one of the most direct physical quantities for detecting MBS. To illustrate the conventional MBS that emerges when spin-rotation symmetry is broken by an external field, we utilise the single FM-AFM $J_{1}$-$J_{2}$ chain. This analysis elucidates the key properties of MBSs, including the magnon binding energy as a function of the ratio between $J_{1}$ and $J_{2}$. These descriptions provide essential information for understanding the internal-field-induced MBSs discussed later.
@@ 2.1 Magnon binding
Magnons are collective excitations that arise from quantised spin wave perturbations in magnetically ordered systems, including ferromagnets, antiferromagnets and helimagnets. These entities adhere to Bose-Einstein statistics, classifying them as bosons. The presence of a magnon signifies a flip in the spin orientation of an electron relative to its ordered (ground) state. It is notable that in environments with strong magnon-magnon interactions and enhanced magnetic correlations, which are typically observed in low-dimensional systems prone to magnetic frustration, magnons exhibit unique behaviours, including the ability to pair and form MBSs. Such pairing results in a composite excitation known as a bimagnon, analogous to Cooper pairs in superconductors.
The quantum phase referred to as a spin nematic state, i.e., a two-magnon bound state, is characterised by a predominance of collective bimagnon excitations. It features a preferred orientation of the paired magnons without spatial confinement, breaking spin-rotational symmetry while retaining translational symmetry. This phenomenon is typically realised in conditions where spins are polarized, for instance, by an external magnetic field. Furthermore, under specific conditions, higher-order bound states involving more than two magnons, such as ‘triatic’ or ‘quartic’ states, may form. These states arise from the intricate interactions of three or four magnons, respectively.
An effective theoretical approach to identifying MBS is to calculate the binding energy of magnons, which is crucial for understanding their interactions and properties. This energy quantifies the cost of dissociating a bound state into individual magnons. This method is analogous to the binding strength of Cooper pairs in superconductivity [27]. When a MBS forms, it is associated with a specific binding energy. To calculate this for a $p$-MBS (where $p$ is the number of magnons, e.g. 2 for nematic, 3 for triatic, 4 for quartic state) at a given net magnetisation $M$ (assuming conserved $S^{z}$), the formula is
@@
where the ground state energy for a total spin $S^{z}=S$ is denoted by $E_{0}(S^{z}=S)$. A positive value of the binding energy in the bulk limit indicates the formation of a bound state, which is an important indicator of the presence of MBS. However, unlike the electron binding energy in superconductivity, the number of bound magnons is not known in advance, so it is necessary to calculate this quantity for different values of $p$. If the largest value of $E_{b}(M,p)$ ($>0$) is given by $p=p_{\rm max}$, it can be concluded that the $p_{\rm max}$ MBS is more energetically favourable than both the individual magnons and other $p$ MBSs. A schematic illustration of the nematic case ($p=2$) is shown in Fig. 2a. The magnon binding energy is observed in experiments as a kind of excitation gap. The following subsection illustrates the calculation of the magnon binding energy for the FM-AFM $J_{1}$-$J_{2}$ chain.
@@
Figure 2: (a) Schematic illustrations of (top) the fully polarized state, (middle) the state with two flipped spins far apart – individual magnons –, and (bottom) the state with two adjacent flipped spins – bounded magnons –. (b) Propagation number $\theta$ (black circles) of the $J_{1}$-$J_{2}$ chain at zero magnetic field as a function of $J_{2}/\lvert J_{1}\rvert$, and the relationship between the number of bound magnons $p$ in the MES at the saturation field. The propagation number is fitted by $\theta/\pi=0.69(J_{2}/\lvert J_{1}\rvert)^{0.29}$ near $J_{2}/\lvert J_{1}\rvert=1/4$. (c) Magnon binding energy $E_{b}(M_{\rm s},p)$ corresponding to $p$ in the MES at the saturation field. The dashed line denotes $E_{b}(M_{\rm s},p)\propto(J_{2}/\lvert J_{1}\rvert)^{\pi/2}$. (d) Ground state phase diagram for MBS as a function of $J_{1}/J_{2}$ (and $J_{2}/\lvert J_{1}\rvert$) and $M/M_{\rm s}$, with the magnon binding energy $E_{b}(M,p)$ depicted in a colour density plot corresponding to each phase.
One of the most promising systems exhibiting MBS is the 1D Heisenberg chain with frustrated FM nearest neighbour (NN) and AFM next nearest neighbour (NNN) interactions, namely the FM-AFM $J_{1}$-$J_{2}$ chain, under an external magnetic field. The Hamiltonian for this system is expressed as follows
@@
The ground state of the system (2) at zero field ($h=0$) has been studied extensively [28, 29, 30, 31, 32, 33, 34]. There are two distinct phases: the FM phase at $J_{2}/\lvert J_{1}\rvert<1/4$ and the incommensurate spiral phase at $J_{2}/\lvert J_{1}\rvert>1/4$. Throughout the spiral phase, the incommensurate spin-spin correlations are short-range [30, 32]; instead, a spontaneous NN-FM dimerisation order accompanied by translational symmetry breaking occurs due to the order-by-disorder mechanism [33, 34]. Consequently, the system exhibits a form of valence bond solid (VBS) state with spin singlet formations between third-neighbour sites [34]. By considering the ferromagnetically dimerised spin-1/2 pair as a spin-1 site, this VBS state can also be recognised as an Affleck-Kennedy-Lieb-Tasaki [35] or Haldane [36, 37] state with topological order protected by global $Z_{2}\times Z_{2}$ symmetry [38].
The theoretical phase diagram at finite magnetic fields has attracted considerable attention in the context of MBSs [10, 11, 12, 13, 14, 15]. When an external magnetic field is applied to the spiral phase, a vector chiral phase appears. This phase is characterised by the breaking of the discrete parity symmetry ($Z_{2}$ symmetry breaking) accompanied by a spontaneous $S^{z}$ spin current circulation [14]. The longitudinal vector chiral order parameter, $\kappa_{l}^{(n)}=\langle(\mbox{\boldmath$S$}_{l}\times\mbox{\boldmath$S$}_{l+n})^{z}\rangle$, acquires a non-zero value in this phase. As the magnetic field strength increases, an MBS appears. The number of bound magnons depends on the ratio of $J_{2}$ to $\lvert J_{1}\rvert$. By calculating the binding energy $E_{\rm b}(M,p)$, expressed in eq. 1, the number of bound magnons can be determined. Fig. 2b shows the propagation number $\theta$ at zero field and the number of bound magnons $p$ near full saturation ($M/M_{\rm s}=1$) as a function of $J_{2}/\lvert J_{1}\rvert-1/4$. The corresponding binding energy is plotted in Fig. 2c. Approaching the FM phase transition at $J_{2}/\lvert J_{1}\rvert=1/4$, the number of bound magnons increases while the binding energy decreases sharply, indicating that larger bound magnon sizes are more susceptible to external perturbations. The conjectured relationship between $\theta$ and $p$, $1/p>\theta/\pi>1/(p+1)$ [15], holds over a wide range of $J_{2}/\lvert J_{1}\rvert$ values, except near $J_{2}/\lvert J_{1}\rvert=1/4$. Furthermore, Fig. 2d shows the distribution of $E_{\rm b}(M,p)$ as a function of $J_{1}/J_{2}$ and $M/M_{\rm s}$. We observe that the binding energy peaks slightly below the saturation magnetisation $M/M_{\rm s}=1$, suggesting that sufficient breaking of the spin-rotation symmetry allows bound magnons to move parallel or antiparallel to the magnetic field, thereby gaining propagation energy.

### v3

@@ Abstract
Multimagnon bound states were predicted nearly a century ago and have since been a key topic in condensed matter physics due to their intriguing quantum properties. However, their realization in natural materials remains elusive, especially in low-dimensional quantum magnets, where stabilizing them is particularly challenging due to the traditionally required extreme external magnetic fields. Therefore, we introduce a novel mechanism that enables the stabilization of multimagnon bound states in quasi-one-dimensional edge-shared cuprates. Our theoretical framework, supported by numerical simulations and experimental data, demonstrates that small antiferromagnetic interchain couplings act as effective internal magnetic fields, promoting a collinear antiferromagnetic order and enabling magnon condensation even at zero external field. This intrinsic stabilisation mechanism eliminates the need for high external fields, offering a platform that is more accessible for experimental realization. We validate this concept by applying it to representative materials such as Li2CuO2, Ca2Y2Cu5555O10, LiCuSbO4, and PbCuSO4(OH)2. Beyond its experimental feasibility, this mechanism could drive advancements in magnon-based quantum computing, low-power spintronic devices, and high-speed magnonic circuits. Moreover, our findings reveal that small interchain and/or interlayer couplings can generally unlock previously overlooked magnetic phenomena, redefining the nature of magnetically ordered states and expanding the frontiers of quantum magnetism.
@@
magnon bound states, Bose-Einstein Condensation, quantum magnetism, spin frustration, quantum phase transitions, low-dimensional spin systems, density-matrix renormalization group
@@ 1 Introduction
Understanding the dynamics of magnons in quantum magnetism is a captivating and evolving area of condensed matter physics. Magnons, the quasiparticles associated with spin waves in magnetically ordered systems, exhibit unique collective behaviors, including the formation of multimagnon bound states (MBSs). Over 90 years ago, Bethe [1] predicted the existence of two-magnon bound states in one-dimensional (1D) quantum magnets, where ferromagnetic (FM) interactions bind two spin excitations together [1, 2, 3]. Since then, the study of MBSs has become central to exploring the interplay between quantum interactions and collective excitations. Beyond the theoretical interest, MBSs have significant implications for quantum computing [4], spintronics [5], and emerging technologies like magnonics [6], where controlled manipulation of magnons is a key objective.
@@
Advancements in quantum simulation have opened new pathways for probing MBSs. For instance, experiments with ultracold atoms in optical lattices have demonstrated magnon excitations and their binding into two-magnon states in quantum Heisenberg chains [7, 8]. Despite this progress, identifying MBSs in naturally occurring materials remains a critical goal. Pursuing this possibility, theoretical studies [9, 10, 11, 12, 13, 14, 15, 16] have extensively investigated the emergence of MBSs in frustrated ferromagnetic-antiferromagnetic (FM-AFM) $J_{1}$-$J_{2}$ chain systems, which are typical magnetic models for quasi-1D (Q1D) edge-shared cuprates [17]. These studies have demonstrated that MBSs can be induced by a magnetic field, and interestingly, near saturation, the types of MBSs vary with the ratio of $J_{1}$ to $J_{2}$, transitioning among two-magnon (nematic), three-magnon (triatic), four-magnon (quartic), and so on. Subsequent experimental efforts have aimed to identify MBSs, focusing on highly 1D materials such as LiVCuO4, where signatures of nematic states have been observed [18, 19, 20, 21, 22, 23]. Nevertheless, the stabilization of MBSs generally requires high external magnetic fields, making experimental observation challenging due to their fragility against thermal fluctuations and external perturbations [24, 25]. This challenge motivates the search for alternative stabilization mechanisms.
In this study, we propose a novel mechanism for stabilizing MBSs in Q1D edge-shared cuprates, a material class known for its complex magnetic interactions. Our approach departs from conventional reliance on external magnetic fields. Instead, we demonstrate that small AFM interchain couplings can act as effective internal fields, inducing collinear antiferromagnetic order (CAFO) within individual chains. This order stabilizes robust MBSs through a self-organised mechanism. Unlike field-driven MBSs, which are often experimentally challenging to observe, this mechanism relies on intrinsic interchain interactions that are more accessible in real materials. This self-organised magnon condensation offers a promising new pathway for experimental realisation, shifting the paradigm of MBS stabilization. A conceptual picture of our MBS mechanism is illustrated in Fig. 1a.
The implications of this mechanism extend beyond MBSs in Q1D systems. Our results suggest that small interchain couplings can induce previously unobserved magnetic phenomena, providing fresh insights into the nature of magnetically ordered states. This perspective invites a reevaluation of magnetic systems across different dimensionalities, where subtle coupling interactions may unlock hidden magnetic behaviors. By demonstrating that interchain coupling-driven effects can mimic external fields, our findings open new avenues for exploring the interplay between quantum magnetism and material geometry.
In doing so, this work not only advances the understanding of MBSs in Q1D systems but also establishes a foundation for uncovering hidden magnetic behaviors in higher-dimensional systems. We begin by outlining the theoretical framework for MBS formation, emphasizing the role of AFM interchain couplings. Subsequently, we present detailed numerical results and experimental comparisons. We conclude with a discussion on broader implications and future research directions.
@@ 2 Magnon bound states by external field
Figure 1: (a) Schematic illustration of the MBS driven by interchain coupling, corresponding to a triatic state with three magnons bound together. The yellow lines represent transverse spin fluctuations. Crystal structures of (b) Li2CuO2, showing two CuO2 chains per unit cell along the b axis, and (c) Ca2Y2Cu5555O10, showing two CuO2 chains per unit cell along the a axis. Red and blue arrows in (b) and (c) indicate an example of symmetry-broken magnetisation directions in the CAFO state. Structure of interchain couplings for compounds such as (d) Li2CuO2 and Ca2Y2Cu5555O10, and (e) LiCuSbO4 and PbCuSO4(OH)2. (f) Schematic of the chain positions used in the numerical 2-chain, 4-chain, and 8-chain calculations, where the dashed lines denote the interchain coupling (see also Ref. [26]).
Prior to discussing our novel emergence mechanism of MBSs, it is necessary to outline the concept of MBSs and to explain how to theoretically calculate the magnon binding energy, which is one of the most direct physical quantities for detecting MBS. To illustrate the conventional MBS that emerges when spin-rotation symmetry is broken by an external field, we utilize the single FM-AFM $J_{1}$-$J_{2}$ chain. This analysis elucidates the key properties of MBSs, including the magnon binding energy as a function of the ratio between $J_{1}$ and $J_{2}$. These descriptions provide essential information for understanding the internal-field-induced MBSs discussed later.
@@ 2.1 Magnon binding
Magnons are collective excitations that arise from quantised spin wave perturbations in magnetically ordered systems, including ferromagnets, antiferromagnets and helimagnets. They follow Bose-Einstein statistics, classifying them as bosons. The presence of a magnon signifies a flip in the spin orientation of an electron relative to its ordered (ground) state. It is notable that in environments with strong magnon-magnon interactions and enhanced magnetic correlations, which are typically observed in low-dimensional systems prone to magnetic frustration, magnons exhibit unique behaviors, including the ability to pair and form MBSs. Such pairing results in a composite excitation known as a bimagnon, analogous to Cooper pairs in superconductors.
The quantum phase referred to as a spin nematic state, i.e., a two-magnon bound state, is characterised by a predominance of collective bimagnon excitations. It features a preferred orientation of the paired magnons without spatial confinement, breaking spin-rotational symmetry while retaining translational symmetry. This phenomenon is typically realised in conditions where spins are polarised, for instance, by an external magnetic field. Furthermore, under specific conditions, higher-order bound states involving more than two magnons, such as ‘triatic’ or ‘quartic’ states, may form. These states arise from the intricate interactions of three or four magnons, respectively.
An effective theoretical approach to identifying MBS is to calculate the binding energy of magnons, which is crucial for understanding their interactions and properties. This energy quantifies the cost of dissociating a bound state into individual magnons. This method is in analogy with the binding energy of Cooper pairs in superconductivity [27]. When a MBS forms, it is associated with a specific binding energy. To calculate this for a $p$-MBS (where $p$ is the number of magnons, e.g. 2 for nematic, 3 for triatic, 4 for quartic state) at a given net magnetisation $M$ (assuming conserved $S^{z}$), the formula is
@@
where the ground state energy for a total spin $S^{z}=S$ is denoted by $E_{0}(S^{z}=S)$. A positive value of the binding energy in the bulk limit indicates the formation of a bound state, which is an important indicator of the presence of MBS. However, unlike for the electron binding energy in superconductivity, the number of bound magnons is not known in advance, so it is necessary to calculate this quantity for different values of $p$. If the largest value of $E_{b}(M,p)$ ($>0$) is given by $p=p_{\rm max}$, it can be concluded that the $p_{\rm max}$ MBS is more energetically favorable than both the individual magnons and other $p$ MBSs. A schematic illustration of the nematic case ($p=2$) is shown in Fig. 2a. The magnon binding energy is observed in experiments as a kind of excitation gap. The following subsection illustrates the calculation of the magnon binding energy for the FM-AFM $J_{1}$-$J_{2}$ chain.
@@
Figure 2: (a) Schematic illustrations of (top) the fully polarised state, (middle) the state with two flipped spins far apart – individual magnons –, and (bottom) the state with two adjacent flipped spins – bounded magnons –. (b) Propagation number $\theta$ (black circles) of the $J_{1}$-$J_{2}$ chain at zero magnetic field as a function of $J_{2}/\lvert J_{1}\rvert$, and the relationship between the number of bound magnons $p$ in the MES at the saturation field. The propagation number is fitted by $\theta/\pi=0.69(J_{2}/\lvert J_{1}\rvert)^{0.29}$ near $J_{2}/\lvert J_{1}\rvert=1/4$. (c) Magnon binding energy $E_{b}(M_{\rm s},p)$ corresponding to $p$ in the MES at the saturation field. The dashed line denotes $E_{b}(M_{\rm s},p)\propto(J_{2}/\lvert J_{1}\rvert)^{\pi/2}$. (d) Ground state phase diagram for MBS as a function of $J_{1}/J_{2}$ (and $J_{2}/\lvert J_{1}\rvert$) and $M/M_{\rm s}$, with the magnon binding energy $E_{b}(M,p)$ depicted in a colour density plot corresponding to each phase.
One of the most promising systems exhibiting MBS is the 1D Heisenberg chain with frustrated FM nearest neighbor (NN) and AFM next nearest neighbor (NNN) interactions, namely the FM-AFM $J_{1}$-$J_{2}$ chain, under an external magnetic field. The Hamiltonian for this system is expressed as follows
@@
The ground state of the system (2) at zero field ($h=0$) has been studied extensively [28, 29, 30, 31, 32, 33, 34]. There are two distinct phases: the FM phase at $J_{2}/\lvert J_{1}\rvert<1/4$ and the incommensurate spiral phase at $J_{2}/\lvert J_{1}\rvert>1/4$. Throughout the spiral phase, the incommensurate spin-spin correlations are short-range [30, 32]; instead, a spontaneous NN-FM dimerization order accompanied by translational symmetry breaking occurs due to the order-by-disorder mechanism [33, 34]. Consequently, the system exhibits a form of valence bond solid (VBS) state with spin singlet formations between third-neighbor sites [34]. By considering the ferromagnetically dimerised spin-1/2 pair as a spin-1 site, this VBS state can also be recognised as an Affleck-Kennedy-Lieb-Tasaki [35] or Haldane [36, 37] state with topological order protected by global $Z_{2}\times Z_{2}$ symmetry [38].
The theoretical phase diagram at finite magnetic fields has attracted considerable attention in the context of MBSs [9, 10, 11, 12, 13, 14]. When an external magnetic field is applied to the spiral phase, a vector chiral phase appears. This phase is characterised by the breaking of the discrete parity symmetry ($Z_{2}$ symmetry breaking) accompanied by a spontaneous $S^{z}$ spin current circulation [13]. The longitudinal vector chiral order parameter, $\kappa_{l}^{(n)}=\langle(\mbox{\boldmath$S$}_{l}\times\mbox{\boldmath$S$}_{l+n})^{z}\rangle$, acquires a non-zero value in this phase. As the magnetic field strength increases, an MBS appears. The number of bound magnons depends on the ratio of $J_{2}$ to $\lvert J_{1}\rvert$. By calculating the binding energy $E_{\rm b}(M,p)$, expressed in eq. 1, the number of bound magnons can be determined. Fig. 2b shows the propagation number $\theta$ at zero field and the number of bound magnons $p$ near full saturation ($M/M_{\rm s}=1$) as a function of $J_{2}/\lvert J_{1}\rvert-1/4$. The corresponding binding energy in the thermodynamic limit is plotted in Fig. 2c. Approaching the FM phase transition at $J_{2}/\lvert J_{1}\rvert=1/4$, the number of bound magnons increases while the binding energy decreases sharply, indicating that larger bound magnon sizes are more susceptible to external perturbations. The conjectured relationship between $\theta$ and $p$, $1/p>\theta/\pi>1/(p+1)$ [14], holds over a wide range of $J_{2}/\lvert J_{1}\rvert$ values, except near $J_{2}/\lvert J_{1}\rvert=1/4$. Furthermore, Fig. 2d shows the distribution of $E_{\rm b}(M,p)$ as a function of $J_{1}/J_{2}$ and $M/M_{\rm s}$. We observe that the binding energy peaks slightly below the saturation magnetisation $M/M_{\rm s}=1$, suggesting that sufficient breaking of the spin-rotation symmetry allows bound magnons to move parallel or antiparallel to the magnetic field, thereby gaining propagation energy.